OpenStudy (anonymous):

graph the function y=3cos(x+pi/2) to draw the graph, plot all x-intercepts, minima, and maxima within one period.

OpenStudy (anonymous):

ok so one period is \[2\pi\] becuase adding \[\frac{\pi}{2}\] doe not change the period

OpenStudy (anonymous):

can you graph it like it asks? please:)

OpenStudy (anonymous):

sure, but first, you should know that the amplitude is 3, because of the 3 out front.

OpenStudy (anonymous):

it might also help to know that \[\cos(x+\frac{\pi}{2})=-\sin(x)\] so this is going to look just like \[-3\sin(x)\]

OpenStudy (anonymous):

if i was going to do this by hand, since the period is \[2\pi\] i would plot the points where x i s \[0,\frac{\pi}{2}, \pi, \frac{3\pi}{2}, 2\pi\]

OpenStudy (anonymous):

if x = 0 you get \[3\cos(0+\frac{\pi}{2})=3\cos(\frac{\pi}{2})=0\] so we know we start at (0, 0) then \[3\cos(\frac{\pi}{2}+\frac{\pi}{2})=3\cos(\pi)=-3\] so we know we are headed down

OpenStudy (anonymous):

then to be brief, back to zero and then up to 3, and then we repeat. it will look like this http://www.wolframalpha.com/input/?i=3cos%28x%2Bpi%2F2%29

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw